A square has sides of length 3x + 1 meters. find a polynomial a(x) that represents the area of the square. find a(1).
Found 2 solutions by Fombitz, mananth:
Answer by Fombitz(32388)
Answer by mananth(16949)

A square has sides of length 3x + 1 meters. find a polynomial a(x) that represents the area of the square. find a(1).
..
length = (3x+1)
Area = L^2
Area = (3x+1)^2
a(x) = 9x^2+6x+1
...
a(1)=9(1)+6(1)+1
a(1)=9+6+1
a(1)=16
...
m.ananth@hotmail.ca